the sum of two numbers is120 if larger number is 4 time smaller number. find the two number?
Answer by ptaylor(2198)

let x=smaller number
then 4x=larger number
Now we are told the following:
x+4x=120 or
5x=120 divide each side by 5
x=24------------the smaller number
4x=4*24=96----------------the larger number
CK
24+96=120
120=120
Does this help?-----ptaylor